### v1

# VHU-Net: Variational Hadamard U-Net for Body MRI Bias Field Correction

Abstract
Bias field artifacts in magnetic resonance imaging (MRI) scans introduce spatially smooth intensity inhomogeneities that degrade image quality and hinder downstream analysis. To address this challenge, we propose a novel variational Hadamard U-Net (VHU-Net) for effective body MRI bias field correction. The encoder comprises multiple convolutional Hadamard transform blocks (ConvHTBlocks), each integrating convolutional layers with a Hadamard transform (HT) layer. Specifically, the HT layer performs channel-wise frequency decomposition to isolate low-frequency components, while a subsequent scaling layer and semi-soft thresholding mechanism suppress redundant high-frequency noise. To compensate for the HT layer’s inability to model inter-channel dependencies, the decoder incorporates an inverse HT-reconstructed transformer block, enabling global, frequency-aware attention for the recovery of spatially consistent bias fields. The stacked decoder ConvHTBlocks further enhance the capacity to reconstruct the underlying ground-truth bias field. Building on the principles of variational inference, we formulate a new evidence lower bound (ELBO) as the training objective, promoting sparsity in the latent space while ensuring accurate bias field estimation. Comprehensive experiments on abdominal and prostate MRI datasets demonstrate the superiority of VHU-Net over existing state-of-the-art methods in terms of intensity uniformity, signal fidelity, and tissue contrast. Moreover, the corrected images yield substantial downstream improvements in segmentation accuracy. Our framework offers computational efficiency, interpretability, and robust performance across multi-center datasets, making it suitable for clinical deployment.
{IEEEkeywords}
Bias field correction, Hadamard transform, U-Net, variational inference, MRI segmentation
Figure 1: (a) Observed abdominal MRI image; (b) Bias-corrected image; (c) Bias field; (d–f) Frequency spectra corresponding to (a–c), respectively.
1 Introduction
[…]
The bias field represents a smooth, low-frequency multiplicative distortion arising from inhomogeneities in the primary magnetic field of magnetic resonance imaging (MRI) systems [1, 2]. As illustrated in Fig. 1 (a), this distortion alters the intensity values in MRI images and produces heterogeneous gray-level distributions for the same tissue. Such variability in tissue intensity obscures fine details critical for accurate clinical interpretation. Consequently, the bias field poses challenges for a range of quantitative analyses, including precise detection, accurate diagnosis, and robust segmentation [3].
To address this challenge, various sophisticated methods have been developed for bias field correction. Contemporary bias field correction methodologies can be classified into two primary categories: traditional methods and neural network-based deep learning methods. Among the most well-known traditional methods, N4ITK [4] builds upon the original N3 [5] algorithm by integrating a robust B-spline approximation and an advanced hierarchical optimization framework. This methodology has demonstrated superior efficacy in bias field correction for brain images, predicated on the assumption of intensity uniformity across distinct brain tissues [6]. However, this assumption does not hold in body MRI, particularly in abdominal and prostate imaging. These regions contain complex anatomical structures and heterogeneous tissues, leading to substantial inter- and intra-organ intensity variations as shown in Fig. 1 (b). As a result, the performance of N4ITK is compromised in these scenarios. Moreover, N4ITK removes the bias field at the expense of significant computational costs, resulting in a low inference speed [7].
In recent years, neural network-based deep learning methods have been widely used in addressing bias field correction. Smiko et al. introduced an implicit training method for MRI bias field removal, bypassing the need for ground-truth data and enabling non-medical data for model training [8]. This approach generalized better across various anatomies and significantly outperformed traditional methods such as N4ITK in both speed and performance. Besides, Sridhara et al. proposed a deep learning approach leveraging an autoencoding framework to estimate bias fields in MRI images  [9]. By correcting intensity inhomogeneities, this method improved tissue classification accuracy, outperforming traditional histogram-based methods when evaluated against ground-truth results. In addition, Chen et al. introduced a novel end-to-end 3D adversarial bias correction network (ABCnet) for infant brain MRI analysis, addressing challenges posed by the dynamic, heterogeneous intensity contrasts caused by nonuniform myelination [6]. The ABCnet is trained alternately to minimize generative and adversarial losses, incorporating a tissue-aware local intensity uniformity term for reducing intensity variation, alongside terms for bias field smoothness and method robustness. It improves accuracy and efficiency in both simulated and real infant brain datasets. However, these approaches have notable limitations: (1) They often require multiple types of annotations, such as bias field maps and tissue segmentations, which are difficult to obtain in body MRI. (2) They typically rely on leveraging the globally smooth and spatially coherent nature of bias fields to facilitate estimation. However, the inherent complexity of body MRI disrupts such global trends. Anatomical regions such as the abdomen and prostate exhibit diverse and heterogeneous tissue properties, leading to pronounced inter- and intra-organ intensity variations that complicate the separation of tissue structures and bias fields.
The extraction of the low-frequency multiplicative field is a fundamental step in MRI bias field correction, as it substantially influences the quality of the corrected images. Effective frequency-domain analysis methods are essential for optimizing this process. Currently, prominent approaches for frequency-domain analysis encompass the discrete Fourier transform [10], discrete cosine transform [11], discrete wavelet transform [12], and Hadamard transform (HT) [13]. These transformations are extensively applied to effectively isolate distinct frequency-domain features, particularly low-frequency components critical for bias field correction. Among these methods, the Hadamard transform stands out for its computational efficiency, as its transformation matrix consists solely of elements valued at $+1$ or $-1$, significantly simplifying implementation [14]. Despite the demonstrated effectiveness of orthogonal transforms in feature extraction tasks, their integration with advanced architectures, such as U-Nets, for MRI bias field correction remains unexplored.
To address the limitations of existing bias field correction methods, this paper proposes a novel frequency-aware variational Hadamard U-Net (VHU-Net) for efficient body MRI bias correction. Unlike conventional deep learning models that infer the bias field by exploiting its global smoothness and spatial coherence in the spatial domain, our method adopts a frequency-domain perspective. Given that the bias field predominantly resides in low-frequency bands, as shown in Fig. 1 (c) and (f), we project MRI data into the Hadamard domain to isolate and extract informative components more effectively. In the encoder module, each input is passed through a stack of convolutional Hadamard transform blocks (ConvHTBlocks), where each block combines standard convolutional operations with the HT layer that isolates frequency components in a channel-wise manner. The entire correction process is regularized via an ELBO-driven loss function that ensures stable and domain-robust representation learning. In the decoder, the bias field is recovered through an inverse HT-reconstructed transformer block (IHTRTB) and a symmetric set of ConvHTBlocks. In addition, a lightweight hypernetwork dynamically modulates the decoder to improve adaptation under varying correction scenarios. The key contributions of this paper are summarized as follows:
  1. 1)
    We introduce a novel ConvHTBlock that combines spatial-domain convolutions with HT-based frequency decomposition. Each block integrates a trainable scaling layer and a semi-soft thresholding function to selectively preserve low-frequency bias-related components while attenuating high-frequency redundancies, thereby enabling compact latent representations.
  2. 2)
    To address the HT layer’s inability to model inter-channel dependencies, we introduce an IHTRTB in the decoder. This module applies a transformer layer to capture global frequency-aware attention patterns and incorporates a semi-soft thresholding layer to further refine the denoising process. This combination enables accurate and spatially coherent bias field reconstruction.
  3. 3)
    A new probabilistic framework based on variational inference is developed, wherein the evidence lower bound (ELBO) acts as the core optimization objective. This formulation enforces sparsity in the latent Hadamard space by aligning the approximate posterior with a Laplacian distribution, thereby improving the robustness and generalizability of the model.
  4. 4)
    We conduct comprehensive evaluations on both simulated and real T2-weighted MRI datasets encompassing abdominal and prostate regions. The proposed VHU-Net consistently surpasses state-of-the-art bias field correction methods, delivering superior intensity homogeneity, enhanced signal fidelity, and notable improvements in downstream segmentation performance. In addition, the model exhibits fast inference speed, ensuring practical applicability in clinical workflows.
2 Related Work
2.1 Bias Field Modeling
A variety of techniques have been developed to tackle intensity nonuniformity in MRI images, with the problem typically formulated as [15, 16]:
$\mathbf{r}=\mathbf{b}\circ\mathbf{f}+\mathbf{n},$ (1)
where $\mathbf{r}$ denotes the observed MRI image, $\mathbf{b}$ is the multiplicative bias field, $\mathbf{f}$ is the underlying inhomogeneity-free image, and $\mathbf{n}$ represents additive noise. The operator $\circ$ denotes the element-wise product. It is commonly modeled using a polynomial basis [6]. In contemporary machine learning paradigms, rather than relying on explicit parametric models for $\mathbf{b}$, the bias field is inferred as the output of a trainable neural network model $\mathbf{b}_{\alpha}$ parameterized by weights $\alpha$. The correction task is formulated as a joint optimization problem, simultaneously estimating the clean image and the bias field by minimizing a composite loss function:
$\mathcal{L}(\alpha,\kappa)=\mathcal{L}_{\text{recon}}(\mathbf{r},\mathbf{b}_{\alpha}\circ\mathbf{f}_{\kappa})+\lambda\mathcal{R}(\mathbf{b}_{\alpha}),$ (2)
where $\mathbf{f}_{\kappa}$ denotes the estimated true image parameterized by $\kappa$, $\mathcal{L}_{\text{recon}}$ enforces fidelity to the observations, $\mathcal{R}$ imposes smoothness regularization on the bias field, and $\lambda$ controls the strength of the regularization. This learning-based formulation enables the model to adaptively capture complex variations in the bias field that may not be well approximated by classical parametric approaches, thus enhancing robustness across diverse acquisition protocols and anatomical regions.
2.2 The Hadamard transform
The two-dimensional Hadamard transform (HT) is a discrete linear transform widely used in signal and image processing due to its computational efficiency and energy compaction properties [17, 18]. It is based on the Hadamard matrix, $\mathbf{H}_{N}$, an orthogonal matrix of size $N\times N$, which is recursively constructed as follows:
$\mathbf{H}_{1}=1,\mathbf{H}_{2}=\begin{bmatrix}1&1\\ 1&-1\end{bmatrix},\mathbf{H}_{N}=\begin{bmatrix}\mathbf{H}_{N/2}&\mathbf{H}_{N/2}\\ \mathbf{H}_{N/2}&-\mathbf{H}_{N/2}\end{bmatrix},$ (3)
where $N=2^{k}$, $k\in\mathbb{N}$. $\mathbf{H}_{1}$ is the base case, and $\mathbf{H}_{N}$ is constructed iteratively from $\mathbf{H}_{N/2}$. For a given input $\mathbf{x}\in\mathbb{R}^{N\times N}$, the 2D HT is computed by applying the HT to both rows and columns of $\mathbf{x}$. The transformation is expressed as:
$\mathbf{Y}=\mathbf{H}_{N}\mathbf{x}\mathbf{H}_{N},$ (4)
where $\mathbf{H}_{N}\in\mathbb{R}^{N\times N}$ is the Hadamard matrix. The resulting matrix $\mathbf{Y}\in\mathbb{R}^{N\times N}$ represents the coefficients of $\mathbf{x}$ in the Hadamard domain. A notable property of this transform is its capacity to concentrate most of the signal’s energy into a small subset of coefficients, primarily located in the upper-left region of $\mathbf{Y}$, which corresponds to the low-frequency components [19]. This property makes the HT effective for low-frequency feature extraction and bias field estimation. To reconstruct original matrix $\mathbf{x}$, the inverse 2D HT is defined as:
$\widetilde{\mathbf{x}}=\frac{1}{N^{2}}\mathbf{H}_{N}\mathbf{Y}\mathbf{H}_{N},$ (5)
where $\widetilde{\mathbf{x}}$ represents the reconstructed matrix. In this study, we utilize the HT to isolate the low-frequency components of the body MRI image, which capture the bias field characteristics.
2.3 Variational inference
Variational inference has emerged as a powerful framework for approximate Bayesian inference, particularly in settings where exact posterior computation is intractable [20]. Specifically, variational inference introduces a family of tractable distributions and seeks the member that minimizes the Kullback–Leibler (KL) divergence to the true posterior [21]. This formulation enables scalable and efficient uncertainty estimation, making it particularly compatible with deep learning models in medical imaging. For instance, by embedding variational inference into neural network architectures, one can learn distributions over latent representations, thereby enabling tasks such as probabilistic segmentation [22], uncertainty quantification [23], and image synthesis [24]. However, despite its demonstrated success in segmentation and related domains [25], the potential of variational inference remains underexplored in broader computational medical imaging tasks, such as MRI bias field correction.
Figure 2: Bias field correction flow.
3 Methodology
3.1 Problem Formation
To enable efficient body MRI bias field correction, we adopt a probabilistic formulation based on latent variable modeling. Let $\mathbf{x}$ denote the observed input image corrupted by intensity inhomogeneity, and ${\mathbf{z}}$ represent the corresponding ground-truth bias field. To learn a deep learning model that most plausibly generates ${\mathbf{z}}$ conditioned on $\mathbf{x}$, we aim to maximize the marginal log-likelihood $\log p({\mathbf{z}}|\mathbf{x})$ during the training stage, given by:
[…]
where $\mathbf{y}$ denotes a latent variable capturing the underlying structure of the image domain. However, this marginalization is generally intractable due to the complexity of the joint distribution $p(\mathbf{z},\mathbf{y}|\mathbf{x})$. To address this, we employ the variational inference framework by introducing an approximate posterior $q(\mathbf{y}|\mathbf{x})$. This distribution is modeled by the encoder network, which infers the latent code $\mathbf{y}$ from the bias-corrupted input $\mathbf{x}$. Next, the marginal log-likelihood is rewritten as follows:
[…]
Applying Jensen’s inequality to the logarithm yields a lower bound $\mathcal{E}$ on the marginal log-likelihood:
$\log p(\mathbf{z}|\mathbf{x})\geq\mathbb{E}_{q(\mathbf{y}|\mathbf{x})}\left[\log\frac{p(\mathbf{z},\mathbf{y}|\mathbf{x})}{q(\mathbf{y}|\mathbf{x})}\right]=\mathcal{E}.$ (8)
We further decompose the joint distribution $p({\mathbf{z}},\mathbf{y}|\mathbf{x})$ using the chain rule:
$p(\mathbf{z},\mathbf{y}|\mathbf{x})=p(\mathbf{z}|\mathbf{y},\mathbf{x})\cdot p(\mathbf{y}|\mathbf{x}),$ (9)
where $p(\mathbf{y}|\mathbf{x})$ is the true posterior. Substituting Eq. (9) into Eq. (8), we obtain the evidence lower bound (ELBO):
$\mathcal{E}$ $=\mathbb{E}_{q(\mathbf{y}|\mathbf{x})}\left[\log p(\mathbf{z}|\mathbf{y},\mathbf{x})\right]+\mathbb{E}_{q(\mathbf{y}\mid\mathbf{x})}\left[\log\frac{p(\mathbf{y}|\mathbf{x})}{q(\mathbf{y}|\mathbf{x})}\right]$
$=\mathbb{E}_{q(\mathbf{y}|\mathbf{x})}\left[\log p(\mathbf{z}|\mathbf{y},\mathbf{x})\right]-\mathcal{D}_{KL}\left(q(\mathbf{y}|\mathbf{x})\,\|\,p(\mathbf{y}|\mathbf{x})\right),$ (10)
where the distribution $p(\mathbf{z}|\mathbf{y},\mathbf{x})$ describes the probability that the decoder reconstructs the ground-truth bias field $\mathbf{z}$ from $\mathbf{y}$ and $\mathbf{x}$. The expected log-likelihood of this distribution quantifies the bias field estimation accuracy. Moreover, the second term represents the negative Kullback–Leibler divergence, which measures the discrepancy between the approximate posterior $q(\mathbf{y}|\mathbf{x})$ and the true posterior $p(\mathbf{y}|\mathbf{x})$.
Unlike the intractable marginal likelihood in Eq. (6), the ELBO is computationally feasible and provides a practical surrogate objective for model optimization. A detailed derivation is provided in Section 3.3. According to Eq. (10), the key to effective bias field estimation lies in maximizing the ELBO. It can be achieved by promoting accurate reconstructions of ground-truth bias field while simultaneously minimizing the divergence between the approximate and true posteriors by enforcing structured regularization on the latent representations.
3.2 Variational Hadamard U-Net (VHU-Net)
To maximize the ELBO, this section presents a novel variational Hadamard U-Net (VHU-Net) that combines the U-Net structure [26] with HT to extract the low-frequency scalar field. It features an encoder-decoder design with symmetric skip connections, allowing for the integration of low-level and high-level features.
3.2.1 Encoder module of the VHU-Net
The encoder of VHU-Net is designed to extract important latent space features through a hybrid framework that integrates spatial-domain convolutions with Hadamard-domain sparsification. As illustrated in Fig. 2, the encoder comprises six hierarchical ConvHTBlocks. Each ConvHTBlock begins with two Visual Geometry Group (VGG) blocks [27] to extract localized features. As depicted in Fig 3, each VGG block comprises a convolutional layer, an instance normalization layer and a leaky ReLU (LReLU) activation function, facilitating stable gradient flow and enhancing nonlinear representational capacity.
To capture bias-related frequency features, we introduce the HT layer as a frequency decomposition module within each ConvHTBlock. Specifically, the HT is employed to transform data $\mathbf{x}\in\mathbb{R}^{C\times H\times W}$ from the spatial domain to the Hadamard domain using Eq. (4), where $C$, $H$, and $W$ denote the number of channels, the height, and the width of the tensor, respectively. Subsequently, we apply three critical steps to accentuate the energy within the low-frequency band, which predominantly contains the bias field information. First, the trainable scaling layer adaptively adjusts the magnitude of frequency components, amplifying the low-frequency signals while diminishing high-frequency noise. It is computed as:
$\widetilde{\mathbf{X}}=\mathbf{\Theta}\circ{\mathbf{X}},$ (11)
where $\mathbf{\Theta}\in\mathbb{R}^{H\times W}$ represents the scaling parameter matrix. ${\mathbf{X}\in\mathbb{R}^{C\times H\times W}}$ denotes the output of the HT. An alternative interpretation of the scaling layer is its relationship with the Hadamard convolution theorem [28, 29], which posits that the element-wise product in the Hadamard domain is analogous to the dyadic convolution in the time domain:
$\mathbf{u}*_{d}{\mathbf{v}}=\mathcal{H}^{-1}\left(\mathcal{H}(\mathbf{u})\circ\mathcal{H}(\mathbf{v})\right),$ (12)
where $*_{d}$ denotes the dyadic convolution. $\mathbf{u}$ and $\mathbf{v}$ represent two input vectors. $\mathcal{H}(\cdot)$ stands for the HT. Hence, the scaling operation in the Hadamard domain can be converted to a convolution in the time domain to enhance feature extraction capability. However, the scaling operation introduces significantly lower computational complexity compared to convolutional filtering in the time domain [30].
Instead of using traditional thresholding techniques, a new trainable semi-soft thresholding layer is developed to eliminate high-frequency components. Unlike hard thresholding, which creates abrupt cutoffs that introduce ringing artifacts [31], or soft thresholding, which uniformly shrinks large coefficients and increases estimation errors [32], our semi-soft approach provides a balanced solution. It completely removes coefficients below the threshold while applying a graduated shrinkage to larger coefficients, maintaining the energy in significant low-frequency components that correspond to the bias field. It is derived from the trainable hard-thresholding operator [33], which is defined as follows:
$\mathcal{Q}({\widetilde{\mathbf{X}},\mathbf{T}})=\mathcal{P}({\widetilde{\mathbf{X}}},\mathbf{T})+\text{sign}(\mathcal{P}({\widetilde{\mathbf{X}}},\mathbf{T}))\circ\mathbf{T},$ (13)
$\mathcal{P}({\widetilde{\mathbf{X}}},\mathbf{T})=\text{sign}({\widetilde{\mathbf{X}}})\circ\text{ReLU}(|{\widetilde{\mathbf{X}}}|-\mathbf{T}),$ (14)
where $\text{ReLU}(\cdot)$ stands for the standard ReLU activation function [34] and $\mathbf{T}\in\mathbb{R}^{H\times W}$ is non-negative trainable threshold parameters. Then, the trainable semi-soft thresholding layer can be computed as follows:
$\mathcal{S}(\widetilde{\mathbf{X}},\mathbf{T})$ $=\text{sign}(\mathcal{Q}({\widetilde{\mathbf{X}},\mathbf{T}})\circ\widetilde{\mathbf{X}})\circ(|\widetilde{\mathbf{X}}|-e^{\mathbf{T}-|\widetilde{\mathbf{X}}|}).$ (15)
The detailed derivation process of Eq. (15) is provided in the supplementary material. After that, we perform the inverse HT (IHT) to transform the image from the Hadamard domain back to the spatial domain using Eq. (5). Through multiple ConvHTBlocks, the VHU-Net processes the input and outputs the latent space representation. Notably, the final ConvHTBlock in the encoder omits the IHT operation as shown in Fig. 2. This design choice preserves the latent representation in the Hadamard domain, allowing the decoder transformer layer to attend to frequency-decomposed features.
Figure 3: (a) ConvHTBlock; (b) HT layer
3.2.2 Decoder module of the VHU-Net
The HT layer performs channel-wise spectral transformation, enabling intra-channel frequency feature extraction. However, it inherently neglects the inter-channel dependencies that are crucial for capturing the global structure of the bias field. To overcome this limitation, an IHT-reconstructed transformer block (IHTRTB) is integrated into the decoder module to implement a global frequency-aware attention mechanism. This design enhances the model’s sensitivity to spatially coherent low-frequency patterns associated with the bias field, while suppressing irrelevant high-frequency information across channels.
Let $\mathbf{Y}\in\mathbb{R}^{S\times M\times N}$ represent the encoded feature map, where $S$ denotes the number of channels and $M\times N$ defines the spatial resolution. The feature tensor is first reshaped into a sequence of spatial tokens $\mathbf{Y}_{\text{seq}}\in\mathbb{R}^{G\times S}$, where $G=M\times N$. This sequence is then passed through two successive transformer blocks. As shown in Fig. 4, each block contains a multi-head self-attention (MHSA) and a multi-layer perceptron (MLP). Both components are preceded by layer normalization and are equipped with residual connections to improve training stability and gradient propagation. In MHSA, the query, key, and value matrices are derived from linear projections:
$\mathbf{Q}_{l}=\mathbf{Y}_{\text{seq}}\mathbf{W}^{Q}_{l},\quad\mathbf{K}_{l}=\mathbf{Y}_{\text{seq}}\mathbf{W}^{K}_{l},\quad\mathbf{V}_{l}=\mathbf{Y}_{\text{seq}}\mathbf{W}^{V}_{l},$ (16)
where $l$ indexes the head number. $\mathbf{W}^{Q}_{l},\mathbf{W}^{K}_{l},\mathbf{W}^{V}_{l}\in\mathbb{R}^{S\times D}$ are learnable projection matrices. $D=\frac{S}{L}$ is the dimensionality of each attention head. $L=8$ is the number of heads. The attention for each head is computed using the scaled dot-product formulation:
$\text{head}_{l}=\text{Attention}(\mathbf{Q}_{l},\mathbf{K}_{l},\mathbf{V}_{l})=\text{Softmax}\left(\frac{\mathbf{Q}_{l}\mathbf{K}_{l}^{\top}}{\sqrt{D}}\right)\mathbf{V}_{l}.$ (17)
Outputs from all $L$ attention heads are concatenated and projected to the original channel dimension:
$\widetilde{\mathbf{Y}}=\text{Concat}(\text{head}_{1},\ldots,\text{head}_{L})\mathbf{W}_{o}+\mathbf{b}_{o}+\mathbf{Y}_{\text{seq}},$ (18)
where $\mathbf{W}_{o}\in\mathbb{R}^{S\times S}$ is the output projection matrix. $\mathbf{b}_{o}\in\mathbb{R}^{S}$ denote trainable bias. Next, the MLP [35] is employed to enhance feature extraction capacity:
$\mathbf{Z}=\text{MLP}(\text{LayerNorm}(\widetilde{\mathbf{Y}}))+\widetilde{\mathbf{Y}},$ (19)
where $\text{LayerNorm}(\cdot)$ refers to layer normalization. After two such transformer blocks, the sequence is reshaped back into its original spatial structure $\mathbf{Z}_{\text{out}}\in\mathbb{R}^{S\times M\times N}$. Then, the semi-soft thresholding layer, as defined in Eq. (15), is applied to remove redundant Hadamard coefficients and introduce nonlinearity after the MLP. The IHT is subsequently performed on the thresholded output to reconstruct the refined bias field.
Figure 4: IHT-reconstructed transformer block.
The decoder mirrors the encoder with five upsampling stages. At each decoding stage, the feature maps from the preceding layer are upsampled and concatenated with the corresponding encoder outputs via skip connections as shown in Fig. 2. The merged feature undergoes a ConvHTBlock.
To enable the decoder to dynamically adapt to different correction conditions, a hypernetwork is introduced to generate channel-wise affine parameters for each decoder ConvHTBlock. This hypernetwork is implemented as a lightweight MLP consisting of three fully connected layers with Leaky ReLU activations. Let $\mathbf{g}^{\ell}\in\mathbb{R}^{C^{\ell}\times H^{\ell}\times W^{\ell}}$ represent the output feature map of the $\ell$-th decoder VGG block, where $C^{\ell}$, $H^{\ell}$, and $W^{\ell}$ correspond to the number of channels, height, and width, respectively. The modulated output $\tilde{\mathbf{g}}^{\ell}$ is computed as:
$\tilde{\mathbf{g}}^{\ell}=\boldsymbol{\gamma}^{\ell}\circ\mathbf{g}^{\ell}+\boldsymbol{\beta}^{\ell},$ (20)
where $\boldsymbol{\gamma}^{\ell},\boldsymbol{\beta}^{\ell}\in\mathbb{R}^{C^{\ell}}$ are the channel-wise scale and bias vectors for the $\ell$-th decoder ConvHTBlock, respectively. All modulation vectors are generated from a shared MLP $\varphi:\mathbb{R}\rightarrow\mathbb{R}^{2\sum_{\ell=0}^{n-1}C_{\ell}}$, conditioned on a scalar control variable $\xi\in\mathbb{R}$:
$[\boldsymbol{\gamma}^{0},\boldsymbol{\beta}^{0},\dots,\boldsymbol{\gamma}^{n-1},\boldsymbol{\beta}^{n-1}]=\varphi(\xi),$ (21)
where $n$ is the number of decoder ConvHTBlocks. Finally, the estimated scalar field is obtained by a $1\times 1$ convolutional layer. The bias-corrected MRI output is then computed by element-wise multiplication between the original input and the estimated scalar field. Here, the scalar field corresponds to the inverse of the bias field.
3.3 The Evidence Lower Bound (ELBO) for VHU-Net
As established in Section 3.1, our objective is to maximize the ELBO to facilitate accurate bias field estimation. This requires jointly minimizing the KL divergence term $\mathcal{D}_{KL}\left(q(\mathbf{y}|\mathbf{x})\,\|\,p(\mathbf{y}|\mathbf{x})\right)$ and maximizing the expected log-likelihood term $\mathbb{E}_{q(\mathbf{y}|\mathbf{x})}\left[\log p({\mathbf{z}}|\mathbf{y},\mathbf{x})\right]$. Within the proposed VHU-Net framework, $q(\mathbf{y}|\mathbf{x})$ stands for the approximate posterior distribution of the latent Hadamard coefficients $\mathbf{y}$, given the bias-corrupted image $\mathbf{x}$. This distribution is learned via the encoder module of VHU-Net.
Prior work on frequency-domain analysis indicates that the HT coefficients, especially the alternating current (AC) components, can be effectively modeled by Laplacian distributions [36]. Additionally, following semi-soft thresholding and filtering, highly correlated components are suppressed. Hence, latent coefficients $y_{i}$ can be modeled as independent and identically distributed Laplacian variables, i.e., $y_{i}\sim La(\mu=0,\omega=b)$ for $11\leq i\leq N-$ 1, where $\mu$ denotes the location parameter, $\omega$ represents the scale parameter and $N$ is the number of HT coefficients in the latent space. Assume $\hat{\mathbf{y}}=[y_{1},\dots,y_{N-1}]$. Accordingly, the probability density function (PDF) of the approximate posterior distribution is given by:
$q(\hat{\mathbf{y}}|\mathbf{x})=\left(\frac{1}{2f_{\omega}(\mathbf{x};\psi)}\right)^{N-1}\exp\left(-\frac{||\hat{\mathbf{y}}||}{f_{\omega}(\mathbf{x};\psi)}\right),$ (22)
where |||||$\cdot$||| | ⋅ | | indicates the 1-norm. $f_{\omega}(\mathbf{x};\psi)$ represents the scale function parameterized by an encoder with learnable weights $\psi$. To better understand the influence of the scale parameter, we analyze the limiting behavior of $q(\hat{\mathbf{y}}|\mathbf{x})$ as $f_{\omega}(\mathbf{x};\psi)\to 0$. For the case where $||\hat{\mathbf{y}}||=0$, we have
$\lim_{f_{\omega}\to 0}q(\hat{\mathbf{y}}|\mathbf{x})=\lim_{f_{\omega}\to 0}\left(\frac{1}{2f_{\omega}(\mathbf{x};\psi)}\right)^{N-1}=\infty.$ (23)
For any fixed $||\hat{\mathbf{y}}||\neq 0$, by applying L’Hôpital’s Rule, we observe:
$\lim_{f_{\omega}\to 0}q(\hat{\mathbf{y}}|\mathbf{x})=\frac{(N-1)!}{(2||\hat{\mathbf{y}}||)^{N-1}}\lim_{f_{\omega}\to 0}\exp\left(-\frac{||\hat{\mathbf{y}}||}{f_{\omega}}\right)=0,$ (24)
where $(N-1)!$ represents the factorial of $N-1$. According to Eq. (23) and Eq. (24), as the value of $f_{\omega}(\mathbf{x};\psi)$ approaches zero, the probability density at $y_{i}=0$ tends toward infinity, while the probability of observing $y_{i}\neq 0$ diminishes to zero. Therefore, setting the value of the scale function $f_{\omega}(\mathbf{x};\psi)$ close to zero can be interpreted as enforcing a sparsity constraint on the Hadamard coefficients in the latent space.
In the context of bias field correction, the bias field predominantly resides in the low-frequency components, whereas high-frequency components are largely redundant. Therefore, for an ideal deep learning model capable of recovering the ground-truth bias field, the latent representation in the Hadamard domain is expected to be sparse. To encode this spectral sparsity in the probabilistic formulation, we assume that the true posterior distribution $p(\hat{\mathbf{y}}|\mathbf{x})$ follows a Laplace distribution with a small scale parameter $\omega$. Its PDF is defined as:
$p(\hat{\mathbf{y}}|\mathbf{x})=(\frac{1}{2\omega})^{N-1}\exp\left(-\frac{||\hat{\mathbf{y}}||}{\omega}\right).$ (25)
The KL divergence between the approximate posterior and true posterior is then given by:
$D_{KL}(q(\hat{\mathbf{y}}|\mathbf{x})\|p(\hat{\mathbf{y}}|\mathbf{x}))$
$=\int q(\hat{\mathbf{y}}|\mathbf{x})\log\frac{q(\hat{\mathbf{y}}|\mathbf{x})}{p(\hat{\mathbf{y}}|\mathbf{x})}\,d\hat{\mathbf{y}}$
$=(N-1)(\frac{f_{\omega}({\mathbf{x}};\psi)}{\omega}+\log\frac{\omega}{f_{\omega}({\mathbf{x}};\psi)}-1).$ (26)
Nevertheless, the specific value of $f_{\omega}({\mathbf{x}};\psi)$ remains undetermined. Subsequently, we employ maximum likelihood estimation to infer the value of scale function $f_{\omega}({\mathbf{x}};\psi)$ from the observed latent AC coefficients $\hat{\mathbf{y}}=[{y}_{1},\ldots,{y}_{N-1}]$. The likelihood of the latent vector $\hat{\mathbf{y}}$ given the estimated scale $f_{\omega}({\mathbf{x}};\psi)$ can be formulated as follows:
${\rm{Lik}}(f_{\omega}({\mathbf{x}};\psi))$ $=p({\hat{\mathbf{y}}}|f_{\omega}({\mathbf{x}};\psi))$
$=\prod_{i=1}^{N-1}\frac{1}{2f_{\omega}({\mathbf{x}};\psi)}\exp\left(-\frac{|{{y}_{i}}|}{f_{\omega}({\mathbf{x}};\psi)}\right).$ (27)
Maximizing the corresponding log-likelihood yields the optimal closed-form estimate for the scale parameter:
$\mathop{\arg\max}\limits_{f_{\omega}({\mathbf{x}};\psi)}\log p({\hat{\mathbf{y}}}|f_{\omega}({\mathbf{x}};\psi))=\frac{1}{N-1}\sum\limits_{i=1}^{N-1}|y_{i}|.$ (28)
Substituting Eq. (28) back into Eq. (26), the KL divergence becomes:
$D_{KL}=(N-1)(\frac{\sum\limits_{i=1}^{N-1}|y_{i}|}{(N-1)\omega}+\log(N-1)\omega-\log{\sum\limits_{i=1}^{N-1}|y_{i}|}-1).$ (29)
It is observed that the minimum of $D_{{KL}}$ is attained when $\sum_{i}|{y}_{i}|=\omega(N-1)$, implying that the latent AC codes are constrained not to zero but to a controlled sparse norm. This constraint regularizes the total latent magnitude to preserve informative frequency components relevant to the bias field, while suppressing high-frequency noise. Additionally, the DC coefficient $y_{0}$ encodes the dominant low-frequency content of the bias field. Enforcing sparsity on $y_{0}$ may result in important information loss. Hence, no regularization is applied to the DC component. Notably, when $\omega$ becomes very small, numerical instability may arise. To address this, the KL divergence can be symmetrically reversed to $D_{\text{KL}}(p(\hat{\mathbf{y}}|\mathbf{x})\|q(\hat{\mathbf{y}}|\mathbf{x}))$, which still encourages sparsity while mitigating overflow risks.
As outlined in Section 3.1, the first term in Eq. (10) represents the likelihood of reconstructing the target scalar field. Maximizing $\mathbb{E}_{q(\mathbf{y}|\mathbf{x})}\left[\log p(\mathbf{z}|\mathbf{y},\mathbf{x})\right]$ can be realized by reducing the mean square error (MSE) between the estimated bias field and the ground-truth map. However, the ground-truth bias field is typically unobservable, rendering direct supervision on the estimated bias field infeasible. To address this limitation, we optimize the network to reduce the MSE between the bias-corrected and the bias-free image. Therefore, the ELBO objective tailored for VHU-Net is formulated as:
$\mathcal{E}_{v}=-\frac{1}{H}\frac{1}{W}\sum_{h=0}^{H-1}\sum_{w=0}^{W-1}\left({e}_{h,w}-{\hat{e}}_{h,w}\right)^{2}-\varepsilon D_{KL},$ (30)
where ${e}_{h,w}$ and $\hat{e}_{h,w}$ are the bias-corrected and bias-free outputs, respectively. $\varepsilon$ is a weighting factor. To further enhance the spatial coherence of the estimated bias field, the final loss function incorporates a smoothness regularization term as defined in Eq. (2):
$\mathcal{L}_{\text{total}}=-\mathcal{E}_{v}+\lambda\mathcal{R}(\mathbf{z}).$ (31)
Here, we employ the Laplacian operator [6] to enforce an explicit smoothness constraint, which penalizes abrupt variations in the predicted bias field.
4 Experiments and Results
4.1 Synthesis of bias field
To simulate realistic intensity inhomogeneities and assess the robustness of correction methods, we evaluated two synthetic bias field generation strategies: (1) Legendre polynomial-based functions [6] and (2) the RandomBiasField function from the TorchIO library [37]. As shown in Fig. 5, Legendre-based fields, constructed using polynomials up to degree $l$, produce smooth, globally varying patterns. However, these tend to lack structural diversity and are often dominated by a single low-frequency mode. In contrast, RandomBiasField introduces spatially heterogeneous bias and yields localized low-frequency variations that more closely resemble the complex bias fields observed in clinical MRI.
We further investigated field complexity by varying the order parameter $r$. As illustrated in Fig. 5, increasing $r$ from 3 to 4 resulted in richer spatial variation and greater irregularity. Compared to $r=3$, the fields generated with $r=4$ exhibit more nuanced structure and variability, better approximating real-world inhomogeneities and posing greater challenges for correction. Accordingly, we adopted RandomBiasField with $r=4$ in all subsequent experiments to ensure realistic and diverse simulation conditions.
Figure 5: Simulation of MRI bias fields.
4.2 Datasets
To evaluate the bias correction performance of the VHU-Net, we performed experiments on two representative categories of body MRI data: abdominal MRI data and prostate MRI data. Additionally, T2-weighted MRI offers superior contrast by emphasizing water-rich tissues, which enhances the clarity of tissue boundaries and facilitates the detection and correction of the bias field. In comparison, T1-weighted MRI exhibits lower tissue contrast and greater signal overlap, particularly with fat, making it more challenging to distinguish bias field from genuine tissue signal variations. For this reason, we focused our experiments on T2-weighted images.
4.2.1 Abdominal MRI data
We collected abdominal MRI scans from the PanSegData dataset [38], comprising 382 real scans and 50 inhomogeneity-free scans, respectively. For model training and evaluation, the real dataset was divided into a training set of 10 scans and a testing set of 372 scans. To address the inherent bias field present in real MRI data, we employed the widely adopted N4ITK algorithm, a standard method in MRI preprocessing over the years. The N4ITK-corrected versions of the 10 training scans served as surrogate ground truth for supervision during training. In addition to evaluating on the real dataset, we created a synthesized dataset for supplementary analysis. This dataset was generated by applying the RandomBiasField function to inhomogeneity-free MRI images. The training set comprised 10 scans, while the testing set included 90 scans.
Table 1: The Qualitative Comparison on Simulated PanSegData
Metrics SSIM ↑ PSNR ↑ COCO ↑ CV $\downarrow$ SNR $\uparrow$ CNR $\uparrow$
GT 1.000 ± 0.000 +$\infty$+ ∞ 1.000 ± 0.000 44.828 ± 0.000 43.974 ± 0.000 9.734 ± 0.000
Input 0.898 ± 0.000 20.632 ± 0.000 0.890 ± 0.000 46.655 ± 0.000 37.223 ± 0.000 4.897 ± 0.000
N4ITK 0.934 ± 0.002 23.817 ± 0.134 0.958 ± 0.002 45.940 ± 0.178 40.895 ± 0.513 5.802 ± 0.278
NPP-Net 0.922 ± 0.010 21.687 ± 0.744 0.923 ± 0.004 45.435 ± 0.186 40.301 ± 0.278 5.059 ± 0.038
CAE 0.945 ± 0.001 25.790 ± 0.306 0.964 ± 0.003 46.371 ± 0.682 41.620 ± 1.023 7.604 ± 0.565
PHU-Net 0.928 ± 0.005 23.250 ± 0.514 0.942 ± 0.008 45.905 ± 0.316 43.813 ± 0.599 4.534 ± 0.185
VHU-Net 0.964 ± 0.000 26.229 ± 0.212 0.987 ± 0.001 44.202 ± 0.061 46.566 ± 0.081 7.788 ± 0.038
Figure 6: The Qualitative Comparison on Simulated PanSegData.
4.2.2 Prostate MRI data
For prostate MRI experiments, we employed four publicly available real datasets: the HK dataset [39], the UCL dataset [39], the RUNMC dataset [40], and the AWS dataset [41]. Similar to the abdominal MRI data preprocessing, the N4ITK was applied to establish the "ground truth". For model training and evaluation, we used a cross-dataset approach to validate generalization performance. All models were trained on the UCL dataset, which provided a training set of 13 scans. Testing was conducted on three other datasets, comprising 12 scans from the HK dataset, 30 scans from the RUNMC dataset, and 48 scans from the AWS dataset.
4.3 Baseline method
To evaluate the bias correction performance of the proposed model, we compare it with N4ITK [4], a well-established standard for bias field correction. Additionally, we extend the analysis by benchmarking the proposed model against other advanced neural network-based deep learning methods, including NPP-Net [26] and CAE [9]. Furthermore, we include a comparison with transform-based learning methods, such as PHU-Net [7], which are particularly effective in extracting low-frequency components like the bias field.
4.4 Metrics for evaluations
To comprehensively evaluate the effectiveness of the proposed bias field correction, we employ a diverse suite of quantitative metrics grouped into three categories. Intensity variation metrics, including the coefficient of variation (CV) [42], contrast-to-noise ratio (CNR) [43], and signal-to-noise ratio (SNR) [44], characterize intra-region uniformity, inter-tissue contrast, and signal clarity, respectively. A lower CV indicates improved tissue homogeneity, while higher CNR and SNR values reflect better contrast and reduced noise, signifying enhanced correction quality. Perceptual similarity metrics assess the consistency between corrected outputs and reference data: the correlation coefficient (COCO) measures the agreement between the estimated and ground-truth bias fields, whereas structural similarity index (SSIM) and peak signal-to-noise ratio (PSNR) evaluate the pixel-level fidelity between the corrected and reference images [16]. For all three, higher values indicate superior performance. Finally, segmentation accuracy metrics, including Dice score, intersection over union (IoU), and positive predictive value (PPV), quantify the overlap and precision of segmentation outcomes [45]. Higher scores in these metrics demonstrate the positive impact of bias correction on downstream anatomical segmentation tasks.
4.5 Results on simulated PanSegData dataset
4.5.1 Quantitative comparison
Table 1 summarizes the quantitative performance of different bias field correction methods on simulated abdominal MRI images. Among all compared methods, the proposed VHU-Net achieves the best overall performance across all metrics. Specifically, it achieves higher SSIM and PSNR than CAE and NPP-Net, indicating better structural preservation and stronger alignment with the true image intensities. The reason is that the limited availability of high-quality phantom data constrains the size of the training dataset. However, CAE and NPP-Net utilize deep convolutional architectures with numerous layers to estimate the bias field, which increases the risk of overfitting. In contrast, VHU-Net incorporates trainable semi-soft-thresholding and the ELBO to enforce sparsity. These sparsity constraints promote the selective activation of a minimal subset of critical neurons, thereby reducing overfitting and enhancing the model’s generalization capability. In terms of COCO, VHU-Net reaches 0.987 ± 0.001, closely approximating the ground-truth bias field and outperforming PHU-Net. This improvement is primarily attributed to the incorporation of a transformer module within the VHU-Net architecture. It enables the model to adaptively attend to informative frequency bands within the transform domain, with particular emphasis on low-frequency regions where bias field components predominantly reside, thereby enhancing estimation accuracy.
Compared to N4ITK, VHU-Net improves intensity homogeneity by reducing the average CV score from 45.940 to 44.202, reflecting a 3.78% enhancement. This improvement stems from the application of a smoothness regularization term during training, which encourages the network to produce spatially smooth bias-corrected outputs by constraining local intensity fluctuations. Additionally, VHU-Net achieves the highest SNR (46.566 ± 0.081) and a significantly improved CNR (7.788 ± 0.038), reflecting enhanced tissue contrast and signal clarity. A key contributing factor to this improvement is the HT layer, which effectively suppresses high-frequency noise in the MRI data. As a result, the model produces denoised outputs with higher SNR than the reference ground truth, highlighting its robustness in noise attenuation. Compared to N4ITK and other deep learning methods, VHU-Net exhibits greater overall consistency and correction quality. These results, coupled with visual assessments, confirm that VHU-Net offers state-of-the-art performance in bias field correction under simulated conditions with known ground truth.
4.5.2 Qualitative comparison
Fig. 6 provides a qualitative comparison of bias field correction methods applied to simulated MRI data. The figure is structured in three rows: the corrected MRI images (top), zoomed-in views for local evaluation (middle), and the estimated bias fields (bottom). All MRI images and bias fields are displayed using pseudo-color mapping to facilitate visual comparison across different methods. In the top row, the corrected MRI images show that VHU-Net achieves superior bias removal, restoring tissue intensities to a visually uniform appearance while preserving anatomical detail. In contrast, NPP-Net leaves noticeable residual inhomogeneities, especially in the right half of the regions. N4ITK and PHU-Net introduce artificial intensity transitions near tissue boundaries. CAE shows improvements over traditional methods but still fails to achieve local consistency comparable to VHU-Net. The second row provides a zoomed-in view of a selected region, enabling finer inspection of local intensity uniformity. VHU-Net exhibits the most consistent intensity distribution within small anatomical structures. Competing methods exhibit varying degrees of local inconsistency, which can negatively impact downstream tasks such as segmentation or quantitative analysis. The bottom row displays the estimated bias fields. Since the ground-truth bias is available in this simulation, qualitative comparison reveals that VHU-Net produces the most accurate and spatially smooth estimation, closely matching the true bias pattern. In comparison, CAE and PHU-Net produce irregular and noisy bias fields, while NPP-Net and N4ITK produce globally smooth fields that fail to capture spatial variations.
Table 2: The Qualitative Comparison on Real PanSegData
Metrics CV $\downarrow$ SNR $\uparrow$ CNR $\uparrow$
Input 47.525 ± 0.000 45.784 ± 0.000 3.574 ± 0.000
N4ITK 46.820 ± 0.016 46.784 ± 0.077 3.806 ± 0.005
NPP-Net 46.219 ± 0.030 48.048 ± 0.042 3.708 ± 0.014
CAE 45.858 ± 1.391 46.645 ± 3.144 3.773 ± 0.112
PHU-Net 46.507 ± 0.063 48.137 ± 0.089 3.686 ± 0.007
VHU-Net 45.308 ± 0.055 48.645 ± 0.013 3.833 ± 0.009
Figure 7: The Qualitative Comparison on Real PanSegData.
4.6 Results on real PanSegData dataset
To comprehensively validate the effectiveness of VHU-Net in real-world clinical scenarios, we applied it to real abdominal MRI images. Unlike the simulated dataset, the real images do not have ground-truth bias fields, which limits the range of evaluable metrics. Therefore, we employed the CV, SNR and CNR as quantitative measures of bias correction performance.
4.6.1 Quantitative comparison
As shown in Table 2, VHU-Net again outperforms all other methods in all reported metrics. It achieves the lowest CV (45.308 ± 0.055), and the highest SNR (48.645 ± 0.013) and CNR (3.833 ± 0.009), indicating superior uniformity, noise suppression, and tissue contrast enhancement. These results demonstrate that VHU-Net not only achieves strong performance on synthetic data but also generalizes effectively to real clinical inputs. Although all neural network-based methods are trained using partially N4ITK-corrected images as reference labels, VHU-Net consistently outperforms N4ITK, CAE, NPP-Net, and PHU-Net on the testing dataset. This improvement can be attributed to fundamental architectural distinctions. Specifically, CAE and NPP-Net utilize deep convolutional architectures with numerous layers to estimate the bias field, which increases the risk of overfitting. In contrast, VHU-Net incorporates the ELBO to enforce sparsity in the Hadamard domain during the training process. The sparsity constraint facilitates the activation of a minimal subset of critical neurons, preventing the model from overfitting to redundant details. This promotes more efficient learning, ultimately improving the model’s generalization ability.
Table 3: The Qualitative Comparison on Real Multi-center Prostate Dataset.
Data Metrics Algorithm
Input N4ITK NPP-Net CAE PHU-Net VHU-Net
HK CV $\downarrow$ 24.285 ± 0.000 24.144 ± 0.100 20.746 ± 0.431 21.430 ± 0.561 20.823 ± 0.003 20.766 ± 0.028
SNR $\uparrow$ 17.851 ± 0.000 15.869 ± 1.401 18.066 ± 0.706 17.582 ± 0.466 19.167 ± 0.093 19.497 ± 0.233
CNR $\uparrow$ 2.012 ± 0.000 1.521 ± 0.347 2.118 ± 0.165 1.996 ± 0.086 2.365 ± 0.022 2.447 ± 0.061
AWS CV $\downarrow$ 37.358 ± 0.000 37.230 ± 0.000 33.407 ± 0.243 33.751 ± 0.292 33.321 ± 0.092 33.108 ± 0.048
SNR $\uparrow$ 26.438 ± 0.000 27.494 ± 0.002 29.577 ± 0.918 26.469 ± 0.304 28.789 ± 0.157 30.460 ± 0.409
CNR $\uparrow$ 3.661 ± 0.000 3.272 ± 0.001 3.788 ± 0.076 3.670 ± 0.032 4.181 ± 0.128 3.914 ± 0.033
RUNMC CV $\downarrow$ 28.346 ± 0.000 28.238 ± 0.076 26.210 ± 0.016 26.553 ± 0.504 26.254 ± 0.090 26.088 ± 0.013
SNR $\uparrow$ 27.944 ± 0.000 28.839 ± 0.595 29.749 ± 0.952 28.124 ± 1.419 29.382 ± 0.239 30.736 ± 0.243
CNR $\uparrow$ 6.373 ± 0.000 6.297 ± 0.058 6.997 ± 0.117 6.532 ± 0.302 7.084 ± 0.147 7.163 ± 0.023
Figure 8: The Qualitative Comparison on Real Multi-center Prostate Dataset.
4.6.2 Qualitative comparison
Fig. 7 illustrates the qualitative performance of various bias field correction methods on the real abdominal dataset. The uncorrected input image exhibits clear low-frequency intensity variations, which compromise tissue contrast and anatomical interpretability. Among the competing methods, N4ITK provides a relatively smooth bias field but fails to adequately correct intensity inhomogeneities near the peripheral regions. Additionally, NPP-Net, CAE and PHU-Net demonstrate strong global correction capability but tend to over-brighten the image, losing fine-grained anatomical features in the process. In contrast, VHU-Net generates a well-structured and coherent bias field, effectively suppressing inhomogeneities while preserving anatomical integrity. The restored image exhibits improved uniformity and structural detail, without introducing over-smoothing or noise artifacts. Notably, the close-up view highlights the superiority of VHU-Net in correcting inhomogeneities. These observations indicate that VHU-Net achieves a better balance between spatial smoothness and correction fidelity, supporting its robustness across diverse anatomical regions.
4.7 Results on real multi-center prostate dataset
4.7.1 Quantitative comparison
To assess the cross-domain generalizability of the proposed method, we conducted experiments on multi-center prostate MRI datasets. Quantitative results are summarized in Table 3. Across all datasets, VHU-Net yields the best performance in terms of both intensity uniformity and contrast enhancement. Specifically, on the HK dataset, it achieves the lowest CV (20.766 ± 0.028) and the highest SNR (19.497 ± 0.233) and CNR (2.447 ± 0.061). Similar trends are observed on the AWS and RUNMC datasets, where VHU-Net outperforms both conventional methods (e.g., N4ITK) and learning-based approaches (e.g., CAE, PHU-Net), demonstrating superior bias field correction and improved image quality. Other deep learning models, while showing moderate gains over the classical baseline, exhibit inconsistencies across different domains. In contrast, VHU-Net maintains high performance regardless of site-specific imaging variations, suggesting strong generalizability and robustness. These results confirm that VHU-Net is not only effective in controlled settings but also scales well across heterogeneous clinical datasets, making it a promising candidate for real-world deployment.
4.7.2 Qualitative comparison
Fig. 8 provides a comparative visualization of bias field correction performance across several representative methods on real prostate MRI images. While NPP-Net generates a generally smooth bias field, it exhibits insufficient compensation near the margins, leaving edge-related artifacts unaddressed. CAE generates a bias field containing high-frequency noise and poor structural alignment due to insufficient spatial regularization. Besides, PHU-Net fails to produce a continuous bias field, instead yielding fragmented and saturated regions that indicate instability in estimation. The proposed VHU-Net addresses these challenges effectively. It produces a bias field that is both smooth and anatomically consistent. The corrected image is visually homogeneous. Additionally, compared to N4ITK, VHU-Net delivers a cleaner and more homogeneous appearance in the close-up area with fewer residual bright patches.
Figure 9: Ablation Experiment Result.
4.8 Ablation study
To validate the effectiveness of each module in the proposed VHU-Net, we performed a comprehensive ablation study on the AWS dataset, as summarized in Fig. 9.
4.8.1 Ablation study on HT layer
When the HT layer is not present in the VHU-Net, the performance degrades, as a higher CV, a lower SNR, and a lower CNR are obtained. It is because the HT decomposes image signals into orthogonal frequency components, allowing the network to isolate and model the low-frequency bias field more effectively. Additionally, the scaling operation in the Hadamard domain is similar to the convolution operation in the spatial domain. This helps extract important features for the bias field estimation.
4.8.2 Ablation study on IHTRTB
Removing the IHTRTB module from VHU-Net leads to a rise in the average CV from 33.108 to 33.216 (↑0.33%), accompanied by reductions in average SNR from 30.460 to 29.232 (↓4.03%) and CNR from 3.914 to 3.821 (↓2.38%), indicating a notable decline in correction performance. The reason is that the IHTRTB employs a frequency-aware attention mechanism that selectively emphasizes informative frequency bands, with a particular focus on low-frequency regions where bias field are concentrated, thereby enabling more accurate bias field estimation.
4.8.3 Ablation study on smoothness regularization
When VHU-Net is trained without smoothness regularization, the model experiences a measurable decline in performance across all evaluation metrics. This is due to the fact that the regularization term plays an essential role in promoting spatial smoothness and penalizing abrupt variations in the predicted bias field, thereby enhancing inhomogeneity suppression of the corrected outputs.
4.8.4 Ablation study on ELBO
To assess the contribution of the ELBO in the proposed framework, we performed an ablation study by excluding the ELBO term from the objective function. As shown in Fig. 9, in the absence of ELBO, the CV increased from 33.108 to 33.174 while the SNR and CNR dropped from 30.460 to 30.230 and 3.914 to 3.832, respectively. By enforcing sparsity and effective regularization in the latent space, the ELBO encourages the elimination of redundant coefficients, such as high-frequency noise and irrelevant details. This facilitates stable optimization and enables more accurate modeling of the underlying bias field.
4.8.5 Ablation study on thresholding layer
Compared to VHU-Net variants incorporating soft-thresholding and hard-thresholding layers, the VHU-Net equipped with semi-soft-thresholding achieves the lowest CV, along with the highest SNR and CNR values. Although the soft-thresholding function removes small high-frequency components, it also attenuates large low-frequency coefficients, leading to errors in bias field estimation. Due to the discontinuity at the threshold, hard-thresholding causes abrupt truncation of frequency components. These sharp spectral transitions introduce discontinuities that manifest in the spatial domain as ringing artifacts [31], ultimately degrading bias field correction quality. However, the semi-soft threshold function eliminates the abrupt discontinuities introduced by the hard-threshold function and alleviates the constant bias issue associated with the soft-threshold function.
5 Conclusion
This study presents VHU-Net, a novel frequency-aware variational architecture designed for effective bias field correction in body MRI. The proposed framework incorporates Hadamard-based ConvHTBlocks in the encoder to isolate low-frequency components associated with the bias field, while suppressing redundant high-frequency noise through a trainable scaling layer and a semi-soft thresholding mechanism. Moreover, the scaling layer applied in the Hadamard domain can be mathematically interpreted as a dyadic convolution operation in the spatial domain. This connection allows VHU-Net to achieve the feature extraction capacity of a convolutional layer while maintaining the computational simplicity of element-wise operations. The decoder integrates an IHTRTB to recover globally coherent and anatomically consistent bias fields by capturing inter-channel relationships. Multiple ConvHTBlocks also enhance the capacity of the decoder to reconstruct accurate bias fields. Additionally, a lightweight hypernetwork dynamically modulates feature responses at each decoding stage via channel-wise affine transformations, enhancing adaptability across anatomical variations. The variational formulation with an ELBO-based objective further ensures sparse, interpretable, and generalizable latent representations. The comprehensive evaluations across simulated and real datasets, including cross-center validation, demonstrate VHU-Net’s robustness and clinical potential. By improving intensity uniformity and tissue contrast, VHU-Net not only enhances visual image quality but also significantly improves downstream segmentation performance, potentially benefiting various clinical applications and quantitative analyses in medical imaging.